An all-equity firm is considering the following projects:

Project Beta IRR

W .58 9%

X .87 9.7%

Y 1.13 12.1%

Z 1.47 15.2%

The T-bill rate is 4.2 percent, and the expected return on the market is 11.2 percent.

Compared with the firm's 11.2 percent cost of capital, Project W has a _____(answer 1) (lower or Higher)____ expected return, Project X has a _____(answer 2) (lower or Higher)____ expected return, Project Y has a _____(answer 3) (lower or Higher)____ expected return, and Project z has a _____(answer 4) (lower or Higher)____ expected return.

Project W should be ___(answer 5) (accepted or rejected)____, Project X should be ___(answer 6) (accepted or rejected)____, Project Y should be ___(answer 7) (accepted or rejected)____, and Project z should be ___(answer 8) (accepted or rejected)____.

If the firm's overall cost of capital were used as a hurdle rate, Project W would be ___(answer 9) (Correctly accepted, Correctly rejected, incorrectly accepted or incorrectly rejected)____, Project X would be ___(answer 10) (Correctly accepted, Correctly rejected, incorrectly accepted or incorrectly rejected)____, Project Y would be ___(answer 11) (Correctly accepted, Correctly rejected, incorrectly accepted or incorrectly rejected)____ and Project Z would be ___(answer 12) (Correctly accepted, Correctly rejected, incorrectly accepted or incorrectly rejected)____
Answer #1

Expected Return of W =Risk free rate+Beta*(Market Return-Risk free Rate)=4.2%+0.58*(11.2%-4.2%) =8.26%
Expected Return of X =Risk free rate+Beta*(Market Return-Risk free Rate)=4.2%+0.87*(11.2%-4.2%) =10.29%
Expected Return of Y =Risk free rate+Beta*(Market Return-Risk free Rate)=4.2%+1.13*(11.2%-4.2%) =12.11%
Expected Return of Z =Risk free rate+Beta*(Market Return-Risk free Rate)=4.2%+1.47*(11.2%-4.2%) =14.49%

If cost of capital is 11.2%
Project W has lower expected return , ​​​​​​​​​​​​​​Project X has lower expected return .​​​​​​​​​​​​​​Project Y has higher expected return .
​​​​​​​​​​​​​​Project Z has higher expected return .

​​​​​​​Project W would be  rejected , ​​​​​​​​​​​​​​​​​​​​​Project X would be  rejected .​​​​​​​​​​​​​​Project Y would be accepted. .
​​​​​​​​​​​​​​Project Z would be accepted.

If IRR would be used as hurdle rate
Project W Correctly rejected.
Project X incorrectly accepted​​​​​​​
Project Y correctly accepted ​​​​​​​
Project Z incorrectly rejected.